### v2

# Process-Supervised Reward Models for Verifying Clinical Note Generation: A Scalable Approach Guided by Domain Expertise

Abstract
Process-supervised reward models (PRMs), which verify large language model (LLM) outputs step-by-step, have achieved significant success in mathematical and coding problems. However, their application to other domains remains largely unexplored. In this work, we train a PRM to provide step-level reward signals for clinical notes generated by LLMs from patient-doctor dialogues. Guided by real-world clinician expertise, we carefully designed step definitions for clinical notes and utilized Gemini-Pro 1.5 to automatically generate process supervision data at scale. Our proposed PRM, trained on the LLaMA-3.1 8B instruct model, outperformed both Gemini-Pro 1.5 and the vanilla outcome-supervised reward model (ORM) in two key evaluations: (1) selecting gold-reference samples from error-containing ones, achieving 98.8% accuracy (versus 70.0% for the vanilla ORM and 93.8% for Gemini-Pro 1.5), and (2) selecting physician-preferred notes, achieving 56.2% accuracy (compared to 37.5% for the vanilla ORM and 50.0% for Gemini-Pro 1.5). Additionally, we conducted ablation studies to determine optimal loss functions and data selection strategies, along with physician reader studies to explore predictors of downstream Best-of-N performance. Our promising results suggest the potential of PRMs to extend beyond the clinical domain, offering a scalable and effective solution for diverse generative tasks.  [^1] Our code is available at https://github.com/hanyin88/prm-clinic.
Process-Supervised Reward Models for Verifying Clinical Note Generation: A Scalable Approach Guided by Domain Expertise
Hanyin Wang1,2, Chufan Gao2, Qiping Xu1, Bolun Liu1, Guleid Hussein1, Hariprasad Korsapati1, Mohamad El Labban1, Kingsley Iheasirim1, , Mohamed Hassan1, Gokhan Anil1, Brian Bartlett1, Jimeng Sun2 1 Mayo Clinic Health System, 2 University of Illinois Urbana-Champaign wang.hanyin@mayo.edu
1 Introduction
Figure 1: Overview of study design. We train a PRM to step-by-step verify clinical notes generated by LLMs from patient-doctor dialogues. Step 1: Clinical notes are transformed into a hierarchical structure of steps, designed based on domain expertise to capture the key considerations in clinical documentation. Step 2: Gemini Pro 1.5 is utilized to generate synthetic errors from predefined categories. These errors are systematically swapped with the original steps to create negative samples. Step 3: The PRM is trained using LLaMA-3.1 8B instruct, followed by ablation studies and physician reader studies to evaluate its performance.
Large Language Models (LLMs) show promising potential in generating clinical notes Van Veen et al. (2024), but their outputs can contain errors and may not always align with physician preferences Omiye et al. (2024); Jin et al. (2024). Currently, no automated, scalable method exists to evaluate the quality of LLM-generated clinical notes, making manual clinician evaluation the gold standard. As a result, costly clinician pilot studies are required to assess LLM-based products, such as ambient AI scribes, in real-world healthcare settings Liu et al. (2024).
A potential solution to this challenge is to verify LLM outputs using a reward model (RM). Reward models are integral to Reinforcement Learning from Human Feedback (RLHF) for LLM post-training Ouyang et al. (2022); Stiennon et al. (2020). When employed as verifiers, RMs are categorized into two types: outcome-supervised reward models (ORMs) Cobbe et al. (2021) and, more recently, process-supervised reward models (PRMs) Lightman et al. (2023). ORMs evaluate an entire generation and assign a single reward score at the end, whereas PRMs provide reward scores at each step of the generation, offering several notable advantages. PRMs enable more fine-grained verification of LLM outputs, improving explainability by pinpointing errors at their exact location. This step-wise reward signaling also facilitates advanced inference-time computation techniques, such as Monte Carlo Tree Search (MCTS), and supports step-by-step reinforcement learning Wang et al. (2023b); Snell et al. (2024).
However, despite significant success in mathematical and coding tasks Lightman et al. (2023); Wang et al. (2023b); Ma et al. (2023), the application of PRMs to other domain remains underexplored Beeching et al. . This gap arises from several key challenges: (1) Verification of correctness: Unlike mathematical or coding problems, which have objective ground-truth answers, verifying correctness in open-ended generative tasks is inherently more complex. (2) Process-supervised data collection: Training PRMs requires datasets with annotations for each step of the answer, which makes it challenging to scale when relying on human annotators.
To address these challenges, we propose an effective method for training PRMs to verify open-ended text generation in the clinical domain. We focus on generating clinical notes from patient-doctor dialogues, an emerging application of LLMs in healthcare Barr et al. (2024). Our approach leverages synthetic data to efficiently generate process-supervised datasets at scale by intentionally introducing errors into high-quality notes. Drawing on domain expertise, we design “errors” and structure “steps” to reflect real-world clinical documentation needs. Our developed PRM, trained on the LLaMA-3.1 8B instruct model, outperformed both Gemini-Pro 1.5 and the vanilla ORM in two key evaluations: (1) selecting gold-reference samples from error-containing ones, achieving 98.8% accuracy (versus 70.0% for the vanilla ORM and 93.8% for Gemini-Pro 1.5), and (2) selecting physician-preferred notes, achieving 56.2% accuracy (compared to 37.5% for the vanilla ORM and 50.0% for Gemini-Pro 1.5). The model also exhibited robust performance in out-of-distribution data. Our main contributions are as follows:
  1. 1.
    PRMs for verifying clinical note generation: We demonstrate that PRMs can effectively provide step-level reward signals for clinical note generation. In real-world deployment, such PRMs serve two practical purposes: (1) automating the evaluation of AI-generated clinical notes to reduce costly clinician reviews, and (2) guiding LLMs to produce higher-quality notes through inference-time computation.
  2. 2.
    Broadening PRM applications beyond math and code: This work expands the application of PRMs beyond their traditional focus on math and coding, demonstrating their potential for open-ended text tasks across diverse domains.
  3. 3.
    Scalable process supervision methodology: We propose a methodology that leverages domain expertise to automatically generate process supervision data at scale.
  4. 4.
    New dataset for the research community: We release our process supervision and physician preference dataset to support future research.
2 Background and Related Work
Verify LLM outputs step-by-step
Recent research highlights a growing interest in process supervision of LLMs to improve their reasoning capabilities. In contrast to earlier work on ORMs Cobbe et al. (2021); Yu et al. (2024), PRMs assess the correctness of LLM outputs step-by-step Lightman et al. (2023); Uesato et al. (2022). One of the core applications of PRMs is Best-of-N sampling, where N complete solutions are generated, and the most suitable one is selected based on the PRM’s reward signal. This method has proven exceptionally effective in mathematical reasoning and code generation tasks, frequently surpassing ORMs Wang et al. (2023b); Ma et al. (2023). Furthermore, PRMs can facilitate step-level reinforcement learning and guide inference-time compute and scaling through techniques such as MCTS Wang et al. (2023b); Ma et al. (2023); Snell et al. (2024). They are considered crucial to the success of advanced reasoning models like OpenAI’s O1. Jiang et al. (2024); Qin et al. (2024); Wang et al. (2024b).
Process-supervised data
Relevant to our work, the need for human-labeled annotations for each step of PRM training data is prohibitively expensive to scale. To address this limitation, several studies have proposed methods to automatically generate process supervision data, including techniques such as Monte Carlo estimation Wang et al. (2023b); Luo et al. (2024); Wang et al. (2024c). However, these approaches are primarily confined to the mathematical domain, where they are effective due to the unambiguous and easily verifiable nature of mathematical solutions. Our exploration of PRM is also related to prior work on evaluating step-by-step reasoning chains Hao et al. (2024), though our focus is on open-ended text generation tasks, which present unique challenges.
Generate clinical note from dialogue
Ambient AI products that transcribe patient-doctor conversations into clinical notes represent one of the most prominent use cases of LLMs in medicine, significantly reducing the burden of clinical documentation Barr et al. (2024). This task was explored at the 2023 ACL ClinicalNLP and CLEFImage workshops, where the best results were achieved using proprietary models Abacha et al. (2023); Yim et al. (2023a). Recent studies have also shown promise in fine-tuning open-source LLMs to generate expert-level clinical notes Wang et al. (2024a, 2023a); Li et al. (2024), though challenges such as inaccuracies and hallucinations remain.
3 Our Approach: Building PRMs to Verify Clinical Note Generation
In this section, we first introduce the problem and task, followed by an explanation of the core method for constructing the process supervision dataset. We then discuss the training, usage, and evaluation methods for PRMs. An overview of our study is illustrated in Figure 1.
3.1 Problem Formation
Given a patient-doctor dialogue, an LLM can generate clinical notes comparable to those written by physicians. We aim to train PRMs to effectively verify LLM-generated notes step-by-step, enabling them to evaluate multiple candidates and select the most accurate and physician-preferred note. Guided by recent work (see details in Appendix D.3), we focus exclusively on the most challenging “Assessment and Plan” (A&P) section of the note.
Figure 2: Two notes to the same dialogue, graded by the PRM. In the top-scoring sample, the note is accurate and concise, with the PRM score (probability of a “+” score for a given step) exceeding 0.5 for each step. In the negative sample, various errors occur across steps, and PRM effectively assigns low scores to those erroneous steps. Items ending in _Score represent the PRM score assigned to each step.
3.2 Training Data Construction
Base Dataset
We utilized the Dialogue-G dataset and modified ACI-BENCH Wang et al. (2024a); Yim et al. (2023b). These datasets comprise synthetic transcripts of patient-doctor dialogues and corresponding notes generated by Gemini Pro 1.0, strictly adhering to the “Best Practice” format recommended by a panel of internal medicine physicians. The original notes from these datasets are considered the gold-reference.
Define Steps
Unlike mathematical problems, where each solution step is typically a single sentence, the concept of a “step” in an open-ended text generation task is less clear. The A&P sections of clinical notes are semi-structured, consisting of multiple problems, each encompassing several clinical reasoning narratives related to their assessment and plan (Figure 1 and 2). While various design options exist, we adopted a heuristic approach informed by physician expertise, as detailed below.
  1. 1.
    The description of each problem is treated as a step, reflecting the importance of the problem list, which is critical for various medical coding and insurance purposes.
  2. 2.
    Within each problem, every sentence constitutes an individual step.
  3. 3.
    After all sentences within each problem, we include a problem-level completeness step. Similarly, after addressing all problems, we add a note-level completeness step. This design aims to prevent the reward model from favoring shorter but incomplete answers.
  4. 4.
    At the end of the note, we introduce an end-of-note step, with its score label representing the overall quality of the note.
Introduce Errors
Through discussions with clinical experts who reviewed LLM-generated notes, we identified three common types of errors:
  • Factual Inaccuracy: Errors involving incorrect information or topics that are referenced in the conversation but not supported by its content.
  • Hallucination: Introduction of entirely unrelated subject entities that were not mentioned in the conversation.
  • Unhelpfulness: Expressions that are vague, incomplete, confusing, or lacking critical details.
We tasked Gemini Pro 1.5 with generating a pool of unique errors for each error type in each case. This was achieved through careful prompt engineering and manual inspection of the errors by our physician co-authors. Subsequently, we randomly swapped original steps in the gold-reference notes with entries from the error pools, ensuring that each error was used only once. At this stage, we introduced a fourth error type, Incompleteness, by randomly removing specific steps or entire problems from the samples.
Introduce Semantic Diversity
Lastly, we tasked Gemini Pro 1.5 with generating a pool of paraphrased sentences to enhance the semantic diversity of the notes. These paraphrases were designed to convey the same information as the original sentences without introducing any new content. We randomly replaced the remaining correct steps in the samples with these paraphrases.
Counts
Total No. of cases by data source
   ACI-BENCH 67
   Dialogue-G 1205
Total No. of samples 9680
Mean No. of samples per case 7.61
Mean No. of errors per sample
   Factual Inaccuracy 1.16
   Hallucination 1.18
   Unhelpfulness 1.19
   Incompleteness 1.27
Mean No. of paraphrases per sample 2.38
Table 1: Summary statistics of PRM-Clinic. The reported numbers are calculated solely from negative samples. Samples from the same case share the same dialogues. During training, gold-reference samples are mixed with negative samples.
Assign Step Scores
Each step is assigned a score label of “$+$” or “--”. The “$+$” label represents a correct step, and the “--” label indicates an incorrect or incomplete step. For gold-reference samples, every step receives a “$+$” label. For samples containing errors, the erroneous step is assigned a score label of “--”. For samples with incompleteness, we adjust the corresponding problem-level or note-level completeness score label to “--”. Lastly, for any samples containing errors, we assign a score label of “--” to the end-of-note step.
PRM-Clinic Dataset
We named our final dataset PRM-Clinic and present its statistics in Table 1. Cases derived from Dialogue-G account for 95.1% of the total instances. For ACI-BENCH, we utilized its training subset Yim et al. (2023b), but with clinical notes generated using Gemini Pro 1.0 Wang et al. (2024a).
Tasks Reference Type Data Source Case Counts Data Distribution Notes
A-Prefer Physician-preference ACI-BENCH tasks 2 and 3 subsets 80 OOD Includes physician preference data used in the first round of RLHF in Wang et al. (2024a). Each case contains three notes generated by an early phase of LLaMA-Clinic.
A-Verify Gold-reference ACI-BENCH tasks 2 and 3 subsets 80 OOD Gold-reference notes generated by an later phase of LLaMA-Clinic. Negative samples are introduced in the same manner as PRM-Clinic.
Dialogue-G Gold-reference Dialogue-G testing subset 80 ID Data constructed in the same manner as PRM-Clinic.
A-Validate Gold-reference ACI-BENCH validation subset 20 ID Data constructed in the same manner as PRM-Clinic.
Table 2: Details of the evaluation tasks. For A-Prefer, A-Verify, and A-Validate, we utilized the original dialogues from the ACI-BENCH dataset. For ID tasks, the notes are generated by Gemini Pro, while for OOD tasks, the notes are generated by LLaMA-Clinic. ID: In-distribution. OOD: Out-of-distribution.
Figure 3: Main results. PRM outperforms both the vanilla ORM and Gemini Pro models across all tasks. Our model can function as either a PRM or ORM. When operating as a PRM, it outperforms our ORM in the A-Prefer task and performs comparably to the ORM in verification tasks. For A-Prefer, the task involves selecting the physician-preferred note from three candidate notes. For all other tasks, the objective is to identify the correct gold-reference sample from negative samples. For ID tasks, the notes are generated by Gemini Pro, while for OOD tasks, the notes are generated by LLaMA-Clinic. The results of the best-performing PRM are presented alongside the ORM results from the same checkpoint. ID: In-distribution. OOD: Out-of-distribution.
3.3 Training of PRMs
The model is trained using a standard pretraining process to optimize the cross-entropy loss:
$L=-\sum_{i\in I}\log p_{\theta}(t_{i}\mid t_{<i}),$
where $\{t_{1},t_{2},\dots,t_{N}\}$ denotes the sequence of tokens in the dataset. The model predicts the probability of the next token given the previous tokens, represented as $p_{\theta}(t_{i}\mid t_{<i})$, with $\theta$ being the model parameters and $t_{<i}=\{t_{1},t_{2},\dots,t_{i-1}\}$. The summation is taken over a set of token positions $I$.
The training corpus consists of dialogue concatenated with clinical notes, where each step in the notes is separated by a step token followed by a score label token. We used LLaMA-3.1 8B Instruct Dubey et al. (2024) as the base model, utilizing its reserved special tokens for all step and score label tokens.
3.4 Usage of PRMs
Use PRMs at Inference Time
At inference time, we perform a single PRM forward pass over the entire clinical note to obtain PRM’s predicted score label for each step. The clinical note itself is generated by a separate LLM. We define the PRM score as the probability that a given step is correct, which is the softmax probability of the special “$+$” token at that step position.
Use Model as PRM or ORM
To compare multiple notes, it is necessary to compute a single note-level score. A key implementation decision is whether to use the reward model as an ORM, which applies when the PRM score at the final end-of-note step serves as the note-level score. In contrast, for PRM, the note-level score is computed as the product of all step-level PRM scores. We experimented with various scoring strategies and found that the product of step-level scores yielded the best results (Appendix C.2).
3.5 Evaluation of PRMs
Evaluation Metrics
We evaluated the reward model’s ability to select the superior note within the Best-of-N framework using two key metrics: (1) Accuracy of selecting the gold-reference note, which measures the model’s ability to identify the gold-reference note among error-containing samples, and (2) Accuracy of selecting the physician-preferred note, which assesses its capability to identify the physician-preferred note from a set of candidate notes.
Notably, the task can be categorized as either in-distribution (ID) or out-of-distribution (OOD). Since our training dataset was developed using the Gemini Pro model family, tasks involving clinical notes generated by other models are considered OOD. For OOD tasks, we used outputs from LLaMA-Clinic. LLaMA-Clinic is a LLaMA-2 13B model that underwent comprehensive domain adaptation for clinical dialogue-to-note generation Wang et al. (2024a). The details of the evaluation dataset are shown in Table 2.
Baseline
We compare our model against three baselines: LLM-as-a-judge using Gemini Pro 1.0 and 1.5, as well as a vanilla ORM implementation. In the vanilla ORM, the training corpus lacks step-level annotations and includes only a single score label for the entire clinical note at the end. For ablation studies, we compare the results of PRMs and ORMs from the same model checkpoint.
Physician Reader Study
To further evaluate the alignment of PRM’s Best-of-N selections with physician preferences, nine physicians reviewed the top selections from various PRM checkpoints across 2,000 generations produced by LLaMA-Clinic. Each experiment involved at least three physicians and was conducted in a randomized and blinded manner.
4 Experimental Results
4.1 Main Results
Figure 3 presents a comparison of model performance across various evaluation tasks, including both ID and ODD scenarios. Across all metrics, our proposed PRM method consistently outperforms the vanilla ORM, as well as both Gemini Pro 1.0 and 1.5. Our model functions as either an ORM or PRM, differing in the aggregation strategy from step-wise to note-level scoring, as detailed in Section 3.4. Notably, our ORM implementation significantly outperforms the vanilla ORM, which predicts only the overall correctness of a generation without learning to estimate step-level reward signals. This suggests that even when functioning as an ORM and relying solely on the final end-of-note score, the ability to predict step-level correctness during training remains essential for achieving strong performance. In verification tasks, PRMs perform comparably to our ORMs when using the same model checkpoint, aligning with prior findings Beeching et al. ; Snell et al. (2024). However, in preference-based evaluations (A-Prefer), PRM surpasses our ORM, achieving 56.2% accuracy compared to 51.2%.
PRM excels in identifying gold-reference samples from error-containing ones in ID datasets, achieving near-perfect accuracy. Therefore, our primary analysis focuses on its generalization performance in OOD settings. Despite the inherent challenges of OOD tasks, PRM exhibits strong generalization, achieving 98.8% accuracy on the A-Verify task. Among all tasks, A-Prefer proves to be the most challenging, reflecting the complexity of physician preferences. Here, PRM achieves 56.2% accuracy in selecting physician-preferred notes among three candidates, significantly outperforming all baselines.
Finally, our results suggest that Gemini Pro models establish a strong baseline when used as LLM-as-a-judge with chain-of-thought reasoning. Notably, Gemini Pro 1.5 exhibits significant improvements over Gemini Pro 1.0 across all metrics, reflecting marked advancements in model capability.
Model Out-of-Distribution In-Distribution A-Prefer A-Verify Dialogue-G A-Validate PRM ORM PRM ORM PRM ORM PRM ORM Baseline (Vanilla Approach)    All Data, Full-Token Loss 46.2 38.8 95.0 97.5 98.8 100.0 100.0 100.0 Ablation: Loss Functions    Score-Token-Only Loss 37.5 27.5 77.5 11.2 81.2 13.8 80.0 10.0    Special-Token Loss 48.8 42.5 90.0 92.5 96.2 95.0 100.0 100.0    Notes-Only Loss 55.0 48.8 91.2 96.2 96.2 98.8 100.0 100.0 Ablation: Data Selection    High Quality Only 50.0 46.2 97.5 97.5 98.8 98.8 100.0 100.0    High Quality Only + Paraphrases 47.5 36.2 95.0 93.8 97.5 96.2 100.0 95.0    High + Medium Quality 45.0 42.5 86.2 97.5 95.0 98.8 100.0 100.0    High + Medium Quality + Paraphrases 43.8 40.0 98.8 100.0 97.5 98.8 100.0 100.0    All Data + paraphrases 43.8 40.0 97.5 98.8 97.5 98.8 100.0 100.0 Notes-Only Loss with Data Selection    High Quality Only 35.0 32.5 87.5 96.2 95.0 98.8 100.0 100.0    High Quality Only + Paraphrases 45.0 38.8 93.8 97.5 97.5 98.8 100.0 100.0    High + Medium Quality 48.8 41.2 93.8 98.8 96.2 96.2 100.0 95.0    High + Medium Quality + Paraphrases 46.2 42.5 96.2 100.0 97.5 98.8 95.0 100.0    All Data + Paraphrases 56.2 51.2 98.8 98.8 98.8 97.5 100.0 100.0
Table 3: Ablation studies on loss functions and data selection. We evaluated different computational strategies for cross-entropy loss and assessed Gemini Pro-filtered data, both with and without paraphrases. The best performing model used notes-only loss and the full dataset with paraphrases. PRM and ORM results are reported from the checkpoint with the highest PRM A-Prefer performance. Numbers represent percentages of accuracy. Bold values indicate the highest PRM and ORM accuracy in each ablation study for ODD tasks. In ID tasks, multiple models achieved perfect accuracy.
4.2 Ablation Studies
In the ablation studies, we examined loss functions and data selection strategies to identify key factors for PRM training. The results are presented in Table 3.
Vanilla Implementation
In the vanilla implementation, the full PRM-Clinic dataset was used without the addition of paraphrases. The training objective is the standard cross-entropy loss applied to all tokens in the corpora, encompassing those from dialogues and notes. While the vanilla PRM model exhibits strong performance on the ID dataset, its performance is weaker on the OOD dataset.
Experiments with Loss Functions
We further investigated the impact of various loss functions, all derived from the same cross-entropy loss framework but differing in the selection of tokens included in the loss computation. The formal definitions of these loss functions are detailed in Appendix A. Restricting the loss computation to only the score label tokens resulted in a marked degradation of performance. Including losses over all special tokens (step tokens and score label tokens) produced performance similar to the vanilla approach but with slightly worse outcomes in verification tasks.
Notably, masking out losses from dialogue tokens while restricting the loss computation to the notes led to a positive impact on performance, particularly in the A-Prefer task. These findings suggest that it is important for the PRM to learn to generate notes (based on the dialogue) while simultaneously learning to predict reward labels in a step-by-step manner. This observation may be influenced by the fact that our PRM did not undergo task-specific supervised fine-tuning prior to reward model training. Masking dialogue losses likely helps prevent overfitting, as dialogues often include lengthy, repetitive content across samples from the same case.
Experiments with Data Selection
While keeping the vanilla loss function, we first evaluated the impact of adding paraphrases to enhance semantic diversity. The results indicate that incorporating paraphrases is beneficial, leading to a general performance improvement in the A-Verify task compared to the baseline, although a slight decrease in A-Prefer accuracy was observed.
During manual inspection of the PRM-Clinic dataset, we identified lower-quality samples in the Dialogue-G subset, characterized by overly brief or unrealistic conversations. To address this, we experimented with selecting only cases with high or medium-quality dialogues, reflecting real-world conversations as judged by Gemini Pro. After filtering out low-quality data, the dataset contains 10,094 samples when retaining high and medium quality, and 5,854 samples when retaining only high quality, compared to the original dataset of 10,952 samples. Despite reducing the total number of training cases by up to 46%, this selection process resulted in modest performance gains.
Best Results
Lastly, we experimented with different data selections while applying cross-entropy loss solely to the notes. Interestingly, the best results came from using the full dataset, which included all quality levels and added paraphrases. For the high-quality data group, masking out the dialogue loss caused a significant drop in performance. This suggests that when the training sample size is small, PRM may depend on learning from dialogue data to enhance step-level reward signal prediction. We used the best-performing PRM for the results shown in Figure 3 and the physician reader study.
Figure 4: Physician Reader Study Results. We selected PRM checkpoints with varying A-Prefer and A-Verify performances for group comparisons. When analyzing the effect of a single metric, we identified two checkpoints with similar scores on the other metric. The Best-of-2000 note from each model was selected for physician review. Each group comparison involved at least three physicians, and the win rate was calculated based on the majority vote.
4.3 Physician Reader Study
In various ablation experiments, we frequently observed a divergence between the trends of A-Prefer (which selects physicians’ preferred samples) and A-Verify (which selects gold-reference samples from those containing errors). To determine which metric better reflects the model’s ability to select the highest-quality notes using Best-of-N in downstream tasks, we conducted a physician reader study. PRM checkpoints were selected based on varying combinations of A-Verify and A-Prefer performance (details provided in Appendix Table 1). The top-ranking notes from their Best-of-2000 selections were then submitted for physician review. As shown in Figure 4, the model with high performance on both metrics notably outperformed the one with low performance on both. Additionally, PRMs with high performance on either A-Prefer or A-Verify demonstrated superior note selection compared to those without, provided the other metric was of comparable value. This finding suggests that both metrics are predictive of Best-of-N performance. Notably, A-Prefer appears to have a greater impact than A-Verify, which is intuitively logical given that A-Prefer directly measures alignment with physician preferences.
5 Conclusions and Future Work
In this work, we proposed a methodology for developing PRMs for open-ended text generation in the clinical domain—guided by domain expertise yet fundamentally simple. We demonstrated the effectiveness of PRM in verifying clinical notes step-by-step and selecting physician-preferred notes via Best-of-N. Informed by key design principles rooted in real-world clinician expertise, our method is scalable and leverages synthetic data to automate process-supervised annotations for PRM training. Notably, our work is among the first to demonstrate the effectiveness of PRMs for open-ended generative tasks beyond mathematics and coding. These promising results suggest the potential for PRMs to extend beyond the clinical domain into a wide range of diverse generative tasks.
While our PRM achieves the highest performance in predicting physician preferences, it is important to note that no current method achieves consistently strong results. This highlights the complexity of the task and suggests that additional factors influencing alignment with physicians remain unexplored, representing valuable opportunities for future research. Lastly, the application of PRM for inference-time compute and scaling (e.g., MCTS) or step-by-step reinforcement learning is an exciting direction for future exploration.
6 Limitations
Our study is constrained by the limited number of physician reviewers and the relatively small dataset involved in the evaluation process. In addition, our approach to the task of clinical note generation builds upon prior work Wang et al. (2024a), particularly their “Best Practice” note format, which was developed based on recommendations from a panel of internal medicine providers. When applying our framework to a different specialty, it may be necessary to recalibrate both the “Best Practice” note format and the corresponding definition of steps for the PRM. Lastly, our work used the proprietary Gemini model to generate synthetic data for PRM training, though the methodology could also be applied to other open-source frameworks (see Appendix D.2 for further discussion).
Appendix A Formal Definitions of Loss Functions for Ablation Studies
A.1 Vanilla Loss
The vanilla implementation computes the cross-entropy loss over all tokens in the dataset, including dialogue and note tokens:
$L_{\text{vanilla}}=-\sum_{i\in I_{\text{all}}}\log p_{\theta}(t_{i}\mid t_{<i}),$
where $I_{\text{all}}$ represents the positions of all tokens in the sequence.
A.2 Loss over Score Label Tokens Only
Focusing on score label tokens only, the loss is computed as:
$L_{\text{score}}=-\sum_{i\in I_{\text{score}}}\log p_{\theta}(t_{i}\mid t_{<i}),$
where $I_{\text{score}}$ denotes the positions of score label tokens.
A.3 Loss over All Special Tokens
When including all special tokens, including both step tokens and score label tokens, the loss is defined as:
$L_{\text{special}}=-\sum_{i\in I_{\text{special}}}\log p_{\theta}(t_{i}\mid t_{<i}),$
with $I_{\text{special}}$ being the set of positions corresponding to special tokens.
A.4 Loss over Notes Only
Restricting the loss to note tokens by masking out dialogue tokens results in:
$L_{\text{notes}}=-\sum_{i\in I_{\text{notes}}}\log p_{\theta}(t_{i}\mid t_{<i}),$
where $I_{\text{notes}}$ represents the positions of note tokens, which also includes all special tokens.
Appendix B Details on Implementation
B.1 Training of PRM
All models were trained using either four 80GB NVIDIA A100 or H100 GPUs, employing DeepSpeed ZeRO Stage 3 optimization and the AdamW optimizer. We adopted Huggingface’s Transformers library and utilized its Trainer module. A limited search was conducted over learning rates [3e-4, 1.5e-4, 3e-5], with 3e-5 selected as the optimal value. The training process utilized consistent hyperparameters across all models, including a global batch size of 16, BF16 precision training, a warm-up phase of 50 steps, and a linear learning rate scheduler. All experiments were trained using one epoch of data.
B.2 Training of Vanilla ORM
For the training of vanilla ORM, we replace the true step score labels in the training corpora with the same placeholder token used during inference, except for the score label at the final end-of-note step. This effectively removes all step-level reward signals, retaining only a single score label for the entire generation during training. Otherwise, we keep the training corpora and pipeline the same as in the rest of the experiments.
B.3 Use PRMs at Inference Time
At inference time, we first used a regular expression to transform the LLM-generated clinical notes into the JSON format defined in our steps. We replace score label tokens with a placeholder token when concatenating the note with the dialogue to form the input context. We then performe a standard forward pass with the input context and obtained the softmax probability of the special “$+$” token at each step position, which represents the PRM score for that step.
When calculating the note-level score, we handle ORM and PRMs differently. For ORM, we simply use the PRM score at the final end-of-note step as the note-level score. For PRMs, to ensure mathematical stability, we avoid directly computing the product of step-level PRM scores. Instead, we take the logarithm of each score and sum the log values across all steps.
B.4 Evaluation Metrics
We calculate accuracy for all metrics at the case level, where each case includes one gold-reference or physician-preferred sample and multiple negative samples. In verification tasks (A-Verify, A-Validation, and Dialogue-G), accuracy is determined by whether the top-scoring sample from the Best-of-N selection, based on PRM or ORM, matches the gold-reference sample. Similarly, in the A-Prefer task, accuracy is determined by whether the top-scoring sample from the Best-of-N selection, based on PRM or ORM, matches the physician-preferred sample.
B.5 Physician Reader Study
The physician reader study was performed using dialogues from ACI-BENCH test1 subset (n = 40). We provided model details for each comparison group in Appendix Table 1. Checkpoints were selected from ablation studies to enable a controlled comparison of the effects of A-Prefer versus A-Verify.
When selecting top-scoring samples for physician review, we first attempt to select those in which the probability of each step’s “$+$” score label exceeds that of its “--” score label. This approach mimics the strategy used in Lightman et al. (2023). If no such samples exist, we select the top-scoring sample from all candidates.
Model (Checkpoint) A-Prefer A-Verify
Dual High vs Dual Low
   All Data + Paraphrases (500) 56.2 98.8
   Score-Token-Only Loss (685) 33.8 86.2
High A-Prefer vs Low A-Prefer
   Notes-Only Loss (685) 55.0 91.2
   Special-Token Loss (685) 43.8 91.2
High A-Verify vs Low A-Verify
   High Quality Only + Paraphrases (366) 45.0 96.2
   High + Medium Quality (631) 45.0 86.2
Appendix Table 1: Details of models used in the physician reader study. Numbers represent percentages of accuracy from PRM.
Of interest, the “Best Practice” note template was recommended by a panel of internal medicine physicians Wang et al. (2024a). We explored collecting preferences from physicians in other specialties, such as obstetrics/gynecology and emergency medicine. As anticipated, there was a high rate of disagreement compared to the preferences of internal medicine physicians, particularly regarding fine-grained aspects such as the perceived target audience (e.g., peer physicians vs. patients) and the desired note length (e.g., concise vs. comprehensive). These findings underscore the complexity of physician preferences and highlight the importance of aligning priorities, criteria, and definitions of “ideal” documentation for real-world deployment. In this work, we report results exclusively from practicing internal medicine physicians.
Appendix C Additional Results
C.1 Distribution of Top-Scoring Samples
Previous studies in the clinical domain have commonly employed lower temperature settings for generative tasks Van Veen et al. (2024); Wang et al. (2024a). In our study, we analyzed the distribution of top-scoring samples derived from PRMs. Specifically, when using LLaMA-Clinic to generate 2,000 samples per case for the physician reader study, we explored various temperature settings. For temperatures in the range of [0.2, 0.4, 0.6, 0.8, and 1.0], we generated 200 samples each. For temperatures ranging from 1.1 to 2.0, samples were generated at 0.1 intervals, with 100 samples per interval.
The distribution of the top-10 scoring samples across all 40 cases is illustrated in Appendix Figure 1. Consistent with prior research, the most frequent temperature range for high-scoring samples was found to be between 0.4 and 1.0.
Appendix Figure 1: Distribution of top-10 scoring samples by PRM across temperature settings. This figure shows the counts of top-scoring samples per temperature setting for the physician reader study, with 2,000 samples generated using LLaMA-Clinic at various temperatures for each case. The highest frequency of top-scoring samples occurred in the temperature range of 0.4 to 1.0.
C.2 Scoring Strategies
We present the A-Prefer results using various scoring strategies for note-level PRM scores, as shown in the Appendix Table 2. These results are derived from the best-performing PRM. The highest performance was achieved by computing the note-level score as the product of all step-level PRM scores.
Product Last Min Mean Median Max Geo Mean Sum Threshold
56.2% 51.2% 45.0% 55.0% 46.2% 42.5% 50.0% 37.5% 38.8%
Appendix Table 2: Accuracy of various scoring strategies for A-Prefer. Each column represents a different method for deriving note-level scores from step-level PRM scores. The Product method multiplies all step-level PRM scores, utilizing the sum of their logarithms during calculations for mathematical stability. The Last method uses the step-level score at the end-of-note token, which is the approach used to obtain scores for ORM. Min takes the minimum score, Mean computes the average, Median uses the median, Max selects the maximum, Geo Mean calculates the geometric mean, Sum adds all scores, and Threshold counts the number of step-level PRM scores that exceed 0.5. Accuracy is determined by whether the top-scoring sample from each method matches the physician-preferred sample. Results are based on the best-performing PRM.
C.3 Quantitative Assessment of In-Distribution and Out-of-Distribution Clinical Notes
To investigate the factors contributing to the robust performance of PRMs in out-of-distribution tasks (A-Prefer and A-Verify), we computed the ROUGE scores of clinical notes generated by physicians, LLaMA-Clinic, and Gemini Pro, based on prior work Wang et al. (2024a) (Appendix Table 3). The high ROUGE scores across various cross-comparison groups suggest strong semantic similarities among all three sources. This result is unsurprising, given that different models should conform to the same best-practice standards and produce similarly high-quality notes for the same case.
Model Comparison ROUGE-1 ROUGE-L ROUGE-Lsum
LLaMA-Clinic vs Physician 0.46 0.34 0.43
Gemini vs Physician 0.49 0.37 0.46
LLaMA-Clinic vs Gemini 0.52 0.40 0.49
Appendix Table 3: Comparison of ROUGE scores among clinical Notes by physicians, LLaMA-Clinic, and Gemini Pro 1.0. Cases are from the physician reader study in Wang et al. (2024a) and include only the “Assessment and Plan” section. Overall, high ROUGE scores were observed for each group comparison.
Appendix D Additional Discussion
D.1 Application of PRMs in Real-World Deployment
Currently, no automated, scalable method exists to evaluate the quality of LLM-generated clinical notes at a fine-grained level. As a result, healthcare institutions rely on costly and time-consuming manual clinician evaluation as the gold standard. Taking the field of ambient AI scribes as an example, a healthcare institution may need to evaluate offerings from multiple industry vendors concurrently. Each evaluation group typically consists of tens to hundreds of clinicians who manually grade LLM-generated clinical notes and provide feedback based on a comprehensive rubric, primarily focusing on detecting errors, inaccuracies, or hallucinations. Furthermore, whenever vendors update their products—for instance, by incorporating a newer version of a foundation model—such evaluations must be repeated.
By applying effectively trained and validated PRMs, the evaluation of LLM-generated notes can be automated at scale, significantly reducing reliance on manual assessment. PRMs enable fine-grained, step-by-step verification of LLM outputs, precisely identifying errors at their exact locations. As demonstrated in our study, PRMs can also select notes that best align with physician preferences. Therefore, PRM-assessed performance can serve as a benchmark for evaluation, akin to the LLM-as-a-judge approach. Furthermore, our study shows that PRMs outperform the vanilla LLM-as-a-judge method. PRMs can also be employed for product monitoring after deployment in healthcare institutions, aiding in the detection of quality issues related to model shift and drift. Importantly, PRMs should first be validated on real-world data and undergo iteration and improvement—for example, through the introduction of additional error types.
Lastly, AI companies can use PRMs to enhance their products. This can be achieved through a simple Best-of-N filtering approach at inference time—similar to our experiments—ensuring that only the highest-quality note is output. Further model improvements are also promising with inference-time scaling techniques such as Monte Carlo Tree Search or step-wise reinforcement learning.
D.2 Training PRMs with an Open-Source Framework
The core of our methodology lies in designing “errors” and structuring “steps” to align with real-world clinical documentation needs. While we opted to use the proprietary Gemini model to generate synthetic errors due to its strong performance and ease of API access, our methodology remains model-agnostic. These synthetic errors could just as effectively be generated using other powerful open-source models, such as LLaMA-3.1 405B Dubey et al. (2024) or Deepseek-R1 Guo et al. (2025) (given adequate GPU resources), provided that the critical decisions regarding error types are made beforehand. Furthermore, our model training is already based on the open-source LLaMA model. Lastly, while we used Gemini Pro to generate synthetic cases for PRM development, PRMs outperform both Gemini Pro 1.0 and 1.5, indicating that the proprietary model does not impose an upper bound on PRM performance within our methodology.
D.3 Focusing on the Assessment and Plan of Clinical Note
Building on recent research on ambient clinical charting and the real-world feedback Wang et al. (2024a), we concentrate solely on the Assessment and Plan (A&P) section of the clinical note generated from patient-doctor dialogues for the following reasons: (1) The A&P section is arguably the most critical and time-intensive part of a clinical note, as it encapsulates key clinical reasoning and medical decision-making. (2) After domain-specific adaptation, the “Subjective” section produced by LLMs is generally of high quality and nearly indistinguishable from human-written content, whereas notable gaps persist in the A&P section. (3) Other sections, such as the “Objective” or “Physical Exam,” are deemed less significant, as they are often imported from the electronic medical record or align less effectively with the physician workflow.
Appendix E Additional Related Work
After our preprint, DeepSeek-R1-Zero and DeepSeek-R1 were released Guo et al. (2025). These models demonstrate remarkable reasoning capabilities through large-scale reinforcement learning with rule-based rewards. In their paper, the authors discussed their reasons for not using PRMs, citing concerns such as reward hacking and the complexity of scaling.
In this exciting new paradigm of large-scale RL, we believe that PRMs remain effective for the following reasons: (1) As a simple yet powerful approach, PRMs serve as an effective means for Best-of-N filtering, guiding test-time computation. (2) In general domains where a verifiable answer is difficult to define, PRMs provide a reliable single reward signal when functioning as ORMs. As demonstrated in our study, this approach outperforms the standard ORM implementation. Moreover, recent research further highlights the potential of PRMs in guiding test-time scaling, offering both effectiveness and computational efficiency Liu et al. (2025); Cheng et al. (2025).
Appendix F Prompts and Instructions
We present the prompts to Geimino Pro for clinical notes-to-JSON format transformation (Appendix Table 4), error generation (Appendix Table 5), paraphrase generation (Appendix Table 6), case quality annotation (Appendix Table 7). and preferred note selection (Appendix Table 8) here. We provided instructions to reviewers in Appendix Table 9.
Appendix G PRM Examples
We presented additional example notes along with their PRM scores in Appendix Table 10 and 11.
Category Prompt
Note to Json Task: Given a clinical note, extract specific information and structure it into a JSON format according to the rules and schema provided.
Instructions:
1. Output Format:
- The response should be in JSON format following the schema outlined below.
2. Sections to Ignore:
- Do not include any content labeled or titled “Assessment and Plan”.
3. Content Extraction:
- Extract the remaining content of the clinical note and organize it into different problems.
- Problems can be identified using a numbered problem list within the note.
- When there’s a section titled “Follow-up instructions:”, treat this as a separate problem.
4. For Each Problem:
- Extract each step as a separate sentence.
- Ignore bullet point symbols, such as -, •, or other similar characters, when extracting steps.
- If the words “Assessment:” or “Plan:” appear in the clinical note, include those words at the beginning of the first sentence that follows their occurrence.
5. Adding Ratings:
- For each step, add a field called “Step_score” with the value “+”.
- For each problem, add a field called “Problem_score” and “Problem_completeness_score” with the value “+”.
6. Add Sequential Numbering:
- For each problem, add a field called “Problem_no”. Number them sequentially starting from “1”, and use strings (e.g., “"1"”, “"2"”).
- For each step, add a field called “Step_no”. Number them sequentially starting from “1”, and use strings.
7. Add Note Score:
- Add a field called “Note_completeness_score” with the value “+” at the root level of the JSON.
8. JSON Schema:
{ “Problems”: [ { “Problem”: “Problem Description”, “Problem_no”: “1”, “Problem_score”: “+”, “Steps”: [ { “Step”: “First step of the problem.”, “Step_no”: “1”, “Step_score”: “+” }, { “Step”: “Second step of the problem.”, “Step_no”: “2”, “Step_score”: “+” } ], “Problem_completeness_score”: “+”, } ], “Note_completeness_score”: “+” }
Here is an example: {Example Note}
Desired output: {Example Json}
Here is the clinical note for your task: {Note}
Appendix Table 4: Prompts to Gemini Pro 1.5 to transform clinical notes into JSON format.
Category Prompt
Prompt Template You are provided with a doctor-patient conversation and its corresponding clinical note in JSON format. Your task is to introduce 10 errors into the clinical note, following the instructions below.
Instructions:
{Error Type Instruction}
- Number of Errors: Introduce 10 errors from the list above at the “Problem” or “Step” level.
- These errors can be introduced at the “Step” field or “Problem” field.
- Do not change other fields such as “Problem_no,” “Problem_score,” “Step_no,” “Step_score” or “Problem_completeness_score.”
- Recording Changes: For each change, only include the following information:
- “Error_type”: The type of error introduced.
- “Problem_no”: The number of the affected problem.
- “Step_no”: The number of the affected step. If change is at the problem level, output null.
- “Error_level”: With a value of “Problem” or “Step.”
- “Detailed_error”: A description of the error introduced.
- “New_content”: The new “Problem” or “Step” content after modification.
- “Original_content”: The original “Problem” or “Step” content before modification.
- Output JSON format with an “Errors” item only, including all information below. Do not include the original JSON file.
Here is the conversation: {dialogue}
Here is the clinical note for your task: {problems}
Factual Inaccuracy Error type is “Factual Inaccuracy”: Introduce detailed factual errors related to the information or topics discussed in the conversation but not supported by it. Examples include changing “left” to “right,” altering medication names, or modifying the follow-up timeframe from “1 month” to “6 months.”
Hallucination Error type is “Hallucination”: Add completely unrelated subject entities that were not discussed in the conversation. This may include fabricated content related to symptoms, diagnostics, treatments, or other aspects. The new information should be major and entirely made up, different from minor factual inaccuracies.
Unhelpfulness Error type is “Unhelpfulness”: Rewrite sentences in a vague, incomplete, or confusing manner. Remove important details, use imprecise language, and avoid specific medical terminology or clear instructions so that the note becomes unhelpful and unclear.
Appendix Table 5: Prompts to Gemini Pro 1.5 for error generation. Each error type is generated using different prompts by incorporating its Error Type Instruction into the prompt template.
Category Prompt
Paraphrase You are provided with a doctor-patient conversation and its corresponding clinical note in JSON format. Your task is to introduce 20 paraphrases based on the original note, following the instructions below.
Instructions:
- You want to paraphrase original sentences to improve semantic diversity of the clinical note. Please make sure the new sentence faithfully represents the same information and knowledge of the original sentence, without adding any new information. Please keep the same academic style but easy to follow, as you would expect from a medical note.
- Number of Errors: Introduce 20 different paraphrases at the “Step” level. Do not introduce paraphrases at the “Problem” level.
- It is ok to introduce multiple paraphrases for the same step. Ideally, we want to introduce paraphrases at various steps.
Here is the conversation: {dialogue}
Here is the clinical note: {problems}
Appendix Table 6: Prompts to Gemini Pro 1.5 for paraphrase generation. Sections of the prompts related to JSON formatting, similar to those in Appendix Table 5, are omitted for brevity.
Category Prompt
Annotate Quality You are provided with a synthetic doctor-patient conversation and its corresponding clinical note in JSON format. Your task is to assess the quality of the conversation and the note.
Instructions:
- Context: The conversations and clinical notes are generated by a large language model. Your task is to assess the quality of these cases by focusing on whether the conversation and notes represent real-world scenarios accurately.
- Conversation Assessment: Evaluate if the conversations are realistic and mimic true clinical interactions during outpatient visits. Be alert for unrealistic conversations, such as interactions involving a newborn speaking or inappropriate dialogue with the mother of a newborn. Identify low-quality conversations that lack sufficient detail or context.
- Clinical Note Assessment: Assess the quality of the clinical note. Identify notes that may contain inaccuracies, hallucinations not supported by the conversation, or that are incoherent or below the standard expected of high-quality medical documentation.
- Recording Changes: Include only the following information in your output.
- Have two root items of “Conversation_quality” and “Note_quality”. For each, include the following items:
- “Rational”: Provide an explanation for your quality assessment.
- “Quality”: Choose among “High,” “Medium,” or “Low.” This indicates your quality assessment.
- “Confidence”: Choose between “High,” “Medium,” or “Low.” This indicates your confidence in the quality assessment.
- Output JSON format according to the specified structure. Do not include the original JSON data in your output.
Here is the conversation: {dialogue}
Here is the clinical note: {problems}
Appendix Table 7: Prompts to Gemini Pro 1.5 for quality annotation. Notably, while we tasked Gemini Pro 1.5 with generating ratings for both conversations and clinical notes, we only used its ratings for conversations in data filtering. This decision was based on manual inspection, which revealed that its ratings and reasoning for clinical notes were of lower quality and accuracy.
Category Prompt
Preferred Note Selection You are given a patient-doctor conversation and several clinical notes based on the conversation. The clinical notes only cover the “Assessment and Plan” section of the note. Your job is to select the best note and provide reasoning.
Here’s the dialogue:
{dialogue}
Here are the notes:
Note 1:
{note_1}
Note 2:
{note_2}
Note 3:
{note_3}
Provide answers in JSON format with two fields: “Rationale” and “Preferred Note”. Explain your reasoning in the “Rationale” field step-by-step. The “Preferred Note” should be the number of the note you select (1, 2, or 3). Make sure your output is in valid JSON format.
Appendix Table 8: Prompts to Gemini Pro for preferred note selection.
Instructions
1. In each row, you will be given a synthetic outpatient patient-provider dialogue from Aci-bench, and two clinical notes based on the same dialogue. We will only evaluate the “Assessment and Plan” parts of a note. 2. We have performed randomization of the notes and simple processing to unify the format of notes. 3. The dialogues from Aci-bench include conversations with (a) calls to a virtual assistant, (b) unconstrained directions or discussions with a scribe, and (c) natural conversations between a doctor and patient. Most conversations occurred in the outpatient setting. 4. Since we are focusing solely on the “Assessment and Plan,” you may assume that all other pertinent information from the dialogue has been documented elsewhere in the note, which is not shown here. Please evaluate the “Assessment and Plan” as you would in a real note. For example, relevant physical exam findings may be helpful in the “Assessment and Plan.” 5. For each row, please start by reading the dialogue and then select your preferred notes. Make your selection based on the overall quality of the note. Essentially, choose the note you would prefer to use in a real patient encounter, imagining you are adopting AI-generated clinical notes for your daily clinical work. You may consider aspects including but not limited to:
(a) Accuracy: Does the information in the clinical note accurately reflect the details from the dialogue?
(b) Completeness: How well does the note cover the important information from the dialogue?
(c) Helpfulness: Does this note include useful information that you would expect from a real “Assessment and Plan”?
However, remember that, ultimately, the selection should reflect your personal preference as a physician.
6. It’s perfectly acceptable to select a tie if you feel that two notes are equally good (or equally poor). 7. Please enter brief comments about each note to help us understand the rationale behind your selection. This is appreciated but not mandatory.
Appendix Table 9: Instructions to physicians for note preference selection.
Category Example
Top-Scoring Sample ASSESSMENT AND PLAN:
1. Allergic Asthma
Assessment: The patient has a history of allergic asthma, recently diagnosed by his primary care physician. He experiences episodic shortness of breath, eye watering, and occasional diarrhea after heavy drinking. Physical exam shows faint expiratory wheezing bilaterally in all lung fields. Chest X-ray and pulmonary function tests were normal.
Plan:
- Continue Albuterol Inhaler as needed.
- Prescribe Singulair 10mg once daily.
- Start allergy testing (skin testing) and try to identify triggers. If unsuccessful, will need further testing in the blood and possibly immunotherapy.
Follow-up instructions:
- Schedule follow-up appointment in one week to review results of skin testing and plan for further treatment.
Negative Sample Problem_1: Newly diagnosed allergic asthma
Step_1: Assessment: Patient reports symptoms of shortness of breath, fatigue, and dry mouth after exposure to his cat.
Step_score: 0.0026
Error: Inaccuracy regarding symptoms after exposure to cat.
Step_4: Differential diagnosis includes COPD, heart failure, pneumonia, bronchitis, PFT abnormalities, vocal cord paralysis, GERD, gastroparesis, esophageal spasm, peptic ulcer disease, irritable bowel syndrome, celiac disease, lactose intolerance, Crohn’s disease, ulcerative colitis, colon cancer, hemorrhoids, anal fissures, fistula, abscess, vasculitis, and sarcoidosis.
Step_score: 0.0086
Error: Hallucination about differential diagnosis.
Problem_2: Lip edema
Problem_score: 0.0420
Error: Hallucination about lip edema.
Step_1: Assessment: Patient has lip swelling and diarrhea following alcohol consumption.
Step_score: 0.2017
Error: In the dialogue the patient answered no to lip edema.
Problem_3: Nausea and vomiting after drinking
Problem_score: 0.0473
Error: Hallucination about these symptoms. In the dialogue the patient answered no to nausea and vomiting.
Appendix Table 10: Example 1 of clinical notes with PRM Scores.
Category Example
Top-Scoring Sample 1. ASSESSMENT AND PLAN:
Hepatitis C
Assessment: The patient’s HCV Ab test was positive, and her liver panel showed an elevated AST (39), but her ALB, Tbili, and Total Bilirubin were all within normal limits. This suggests that she has chronic hepatitis C.
Plan:
- Confirm the diagnosis with additional blood work, including HCV RNA and HCV Genotype.
- Determine the severity of liver disease with a special ultrasound.
- Once the diagnosis and severity are confirmed, discuss treatment options with the patient.
2. Fatigue
Assessment: The patient reports feeling tired at the end of the day after working. She denies any other symptoms.
Plan:
- Recommend the patient to prioritize rest and relaxation.
- Encourage the patient to maintain a healthy sleep schedule.
- Consider referral to a sleep specialist if fatigue persists despite lifestyle modifications.
Follow-up instructions:
- Return to clinic in 3 weeks for follow-up.
Negative Sample Problem_1: Hepatitis C virus (HCV) infection
    Step_7: Plan: - Obtain follow-up blood tests including HCV RNA, HCV genotype, FIBROSIS OF THE LIVER, LFTS, TSH, UA, BMP.
    Step_score: 0.0081
    Error: Inaccuracy about tests ordered.
    Problem_Completeness_score: 0.0159.
    Error: Failed to include liver ultrasound.
Problem_2: Fatigue and mild anemia
    Problem_score: 0.0110
    Error: Hallucination about anemia.
    Step_4: She denies any recent change in bowel movements, urination, fevers, chills, weight loss, joint pain, night sweats, shortness of breath.
    Step_score: 0.2225
    Error: These symptoms are not discussed in the dialogue.
    Step_6: Plan: - Recommend restful activities, stress reduction techniques, counseling on healthy eating habits, exercise.
    Step_score: 0.0293
    Error: These plans are not discussed in the dialogue.
Appendix Table 11: Example 2 of clinical notes with PRM Scores